# An Automated Framework for Large-Scale Graph-Based Cerebrovascular Analysis

### v2

@@ 1 Introduction
Current approaches for cerebrovascular analysis can be classified into three categories. Segmentation-based methods, enabled by recent advances in deep learning for accurate automated vessel segmentation, typically from time-of-flight magnetic resonance angiography (TOF-MRA), estimate arterial volume from the segmented masks at the whole brain and regional levels [2]. Atlas-based approaches provide population references through statistical arterial atlases [3], offering normative cerebrovascular anatomy baselines. However, both approaches focus on simple volumetric and diameter measurements that may fail to completely characterize the vascular tree. Finally, graph-based computational tools use a segmentation mask as input to generate a graph representation of the vascular tree, from which intensity and morphometric features are extracted [4, 5, 6, 7, 8]. While commonly used, these tools are either semi-automatic [4, 7] or have been specifically designed for a particular type of image [5, 6, 8], limiting their usage at scale for any neurovascular imaging modality.
@@
Table 1: List of vascular features extracted in within the vessel analysis framework.
@@
Abnormal degree nodes $|\{v\in V:\deg(v)>3\}|$ Anomalous branching
@@
Fractal Fractal dimension $\frac{\Delta\log(N(\varepsilon))}{\Delta\log(\varepsilon)};$$N(\varepsilon)$ the boxes at scale $\varepsilon$ Multi-scale organization
@@
Mean square curvature $\frac{1}{L_{geo}}\int_{0}^{1}\frac{\kappa^{2}(t)}{[n(t)]^{2}}\|\mathbf{x}^{\prime}(t)\|dt$ Curvature variability
@@
In this work, we present CaravelMetrics, an automated framework for analyzing the cerebrovascular tree. Our framework extracts fifteen morphometric, topological, fractal, and geometric vascular features from a graph generated from vascular binary masks, thus being agnostic to the underlying image modality. Similar to [2], we integrate an atlas of arterial regions, allowing us to analyze the cerebrovascular tree at both the whole-brain and regional levels. We showcase the potential of our open-source tool by analyzing the IXI Dataset [^1] https://brain-development.org/ixi-dataset/, a cohort of 570 subjects aged 20 to 86 years whose cerebrovascular trees were imaged using TOF-MRA. Our analysis reveals that our features effectively capture variations in the vessels across a population.
@@ 2 Methods
CaravelMetrics relies on representing the cerebrovascular network as a mathematical graph structure that is extracted from a vascular binary mask. The graph and the vascular binary mask are then used to extract morphometric, topological, fractal, and geometric features that can be aggregated at the whole-brain or regional level through the usage of an arterial atlas (Figure 1).
Graph Representation of Cerebrovascular Networks. Let us denote $M:\Omega\subset\mathbb{Z}^{3}\to\{0,1\}$, a binary vascular mask obtained from any 3D imaging modality. Given $M$, we define its skeleton $S=\operatorname{Skel}(M)\subseteq\Omega$ as a one-voxel-thick subset that preserves the topological properties of $M$, such that $S\subseteq M$, $\operatorname{conn}(S)=\operatorname{conn}(M)$, and $\operatorname{hom}(S)=\operatorname{hom}(M)$, where $\operatorname{conn}(\cdot)$ denotes the number of connected components and $\operatorname{hom}(\cdot)$ describes the topology (loops and holes).
From the skeleton points $\mathbf{x}=(x,y,z)\in S\subset\mathbb{Z}^{3}$, we construct an undirected weighted graph $G=(V,E)$, where each skeleton point corresponds to a node $v\in V$ with spatial coordinates $\mathbf{x}_{v}\in\mathbb{Z}^{3}$. Edges $E=\{(u,v)\}$ connect pairs of nodes within a $3\times 3\times 3$ neighborhood (26-connectivity), with edge weights defined as the Euclidean distance $d_{uv}=\|u-v\|_{2}$.
@@
Feature Extraction. The framework extracts fifteen complementary vascular features grouped into four categories (Table 1), each capturing a distinct property of cerebrovascular organization. Morphometric features quantify network extent and capacity through total vessel length and volume. Topological features describe branching architecture and connectivity using bifurcation count, loop number, and component analysis. Fractal features characterize hierarchical self-similarity via box-counting and lacunarity [9]. Finally, geometric features quantify vessel tortuosity, reflecting arterial stiffening and age-related alterations, by capturing local vessel morphology and shape irregularity through cubic-spline parameterization of discrete centerline segments (Eq. 1) as continuous curves $\mathbf{x}(t):[0,1]\!\to\!\mathbb{R}^{3}$, with curvature values weighted by path multiplicity $n(t)$ and averaged across root nodes $s\in\{R_{1},R_{2},R_{3}\}$ to obtain descriptors for each $G_{k}$.
@@
The cerebrovascular tree exhibits heterogeneous patterns across brain territories, motivating a region-based analysis across arterial territories. We enable such analysis by aligning an arterial atlas $\mathcal{A}=\{T_{i}\}_{i=1}^{T}$, which partitions the brain into $T$ arterial territories, with the original neurovascular image. This yields territory-specific masks defined as $M_{i}=M\cap T_{i}.$ Our framework supports two complementary analysis modes: whole-brain analysis, performed on the complete graph $G=(V,E)$ derived from $M$; and region-specific analysis, performed on regional graphs $G_{i}=(V_{i},E_{i})$ derived from each $M_{i}$. This formulation enables the extraction of vessel features globally or within anatomical arterial territories.
Implementation Details. Skeletonization, graph generation, and feature extraction are implemented in Python 3.10 with NumPy, SciPy, scikit-image, and NetworkX. We employ the open-source 3D brain MRI arterial territories atlas proposed in [10] ($T=30$). Atlas-to-image alignment is done through multi-step registration: brain extraction via FSL BET [11], reorientation to MNI space, atlas to structural image registration using FLIRT with mutual information, and structural image to neurovascular image co-registration. Code and documentation are available in the project’s repository [^2] https://github.com/i-vesseg/CaravelMetrics.
@@
3 Experiments & Results
We demonstrate the relevance of the proposed framework by analyzing 570 healthy subjects, aged 20–86 years, from the IXI dataset, with multiple MR sequences acquired across three London hospitals with different scanners: Guy’s Hospital (Guys), Hammersmith Hospital (HH), and Institute of Psychiatry (IOP). Each subject has associated demographic information, including age, sex, height, weight, education, and occupation. We compute the body mass index (BMI) from the height and weight, and categorize individuals into 4 BMI-based groups according to standard clinical thresholds.
Using IXI’s TOF-MRA deep learning-generated binary vessel masks from the VesselVerse dataset [12], our aim is to use features from CaravelMetrics to quantify age-related cerebrovascular changes across the healthy adult lifespan and identify demographic factors associated with vessel morphology, tortuosity, and network topology, consistent with the literature (e.g., [13]). For the atlas-based regional analysis, we considered T1w MR images within IXI as structural images for multi-step registration. In our analysis, age associations employ Spearman correlations and one-way ANOVA across quartile-based groups with $\eta^{2}$ effect sizes. Demographic comparisons utilize t-tests for binary factors (sex) and one-way ANOVA for categorical variables (BMI, education), with $\omega^{2}$ effect sizes for unbalanced groups. Multi-center effects undergo site-stratified analysis to separate biological variation from acquisition artifacts. All analyses apply Benjamini-Hochberg False Discovery Rate (FDR) correction ($q<0.05$) to control for multiple comparisons.
Results. Initial analysis revealed substantial site effects, with IOP showing markedly different vessel distributions than Guys and HH (Figure 2), likely reflecting scanner/acquisition variations [3], though age imbalances may have contributed. Excluding IOP and focusing on the two homogeneous sites ($n=488$) strengthened age effects: total vessel length effect size increased from $\eta^{2}=0.137$ to $0.204$ (6.7 percentage points), suggesting site heterogeneity attenuates biological associations.
Analyzing age-related cerebrovascular changes, strong negative correlations emerged between age and structural vessel metrics across the adult lifespan (Figure 3), with total vessel length showing the most pronounced decline (r=0.50r=-0.50, $p<0.001$), corresponding to 20% reduction from youngest to oldest subjects (Figure 4). Volume ($r\approx-0.25$) and bifurcation count ($r\approx-0.25$) also declined significantly, reflecting network simplification. Conversely, bifurcation density and tortuosity increased with age ($r\approx+0.08$ and +0.10+0.10), consistent with arterial stiffening, while fractal dimension and lacunarity showed no significant correlations, indicating preserved topological scaling despite structural changes. Quartile-based age group analysis revealed large effect sizes: total length $\eta^{2}=0.204$, volume $\eta^{2}=0.047$, bifurcations $\eta^{2}=0.064$ (all $p<0.001$). Regional analysis (Figure 5) revealed modest changes, with vessel length showing weak negative correlations in anterior territories and tortuosity increases in posterior and deep brain regions, though correlations remained relatively small across all territories. This pattern likely reflects measurement noise at smaller scales and the systemic nature of cerebrovascular aging, affecting the entire vascular network uniformly.
@@ 3 Experiments & Results
Regarding anthropometric factors, normal-weight individuals ($n=252$) demonstrated longer vessels than overweight ($n=146$) or obese ($n=55$) groups ($\eta^{2}=0.032$, $p=0.007$; Figure 6, left). Height showed positive correlations with multiple vessel metrics ($\eta^{2}=0.026$$0.053$), with taller individuals demonstrating higher bifurcation density, volume, and network complexity, reflecting brain size scaling. Concerning sex, after IOP removal, males ($n=223$) showed higher bifurcation density ($d=0.384$), loops ($d=0.321$), and volume ($d=0.29$), along with lower lacunarity ($d=0.193$), than females ($n=265$), consistent with known neuroanatomical differences [13]. Finally, education analysis was prioritized over occupation given stronger age confounding in the latter (retired individuals averaged 68 years versus 26 for students), revealing a clear stepwise gradient in vessel length across qualification levels ($\eta^{2}=0.055$, $p<0.001$; Figure 6, right), consistent with cognitive reserve extending to cerebrovascular architecture [14].
@@ 4 Discussion and Conclusion
We introduced CaravelMetrics, an automated framework for cerebrovascular characterization that extracts fifteen complementary features encompassing morphometric, topological, fractal, and geometric descriptors. By integrating an arterial atlas, the framework supports both whole-brain and territory-specific analyses, enabling multiscale assessment of cerebrovascular organization. As it operates directly on binary vessel masks, CaravelMetrics is compatible with diverse neurovascular imaging modalities. Applied to 570 subjects from the IXI dataset (ages 20–86), the framework revealed age-related declines in vessel length, volume, and bifurcation count, alongside increased tortuosity and education-associated increases in vessel length. These patterns are consistent with findings reported in the literature, supporting the validity of the proposed framework for cerebrovascular feature extraction and analysis at scale. A key limitation of our approach is its dependence on the quality of the deep learning-generated input masks. While we did not explicitly assess this, results suggest our framework is resilient to moderate segmentation noise. Future research should evaluate this robustness, develop strategies to address inaccuracies, and explore the framework’s applicability in more challenging pathological groups. CaravelMetrics enables reproducible, population-scale analysis of cerebrovascular networks that can help understand brain health and aging.
@@ 6 Acknowledgments
This work is co-funded by the European Union (ERC CARAVEL 101171357), and supported by France Life Imaging (ANR-11-INBS-0006). LSC is supported by TwinsUK Imaging: A Resource for Ageing Research (Chronic Disease Research Foundation). MM and SO are supported by King’s Health Partners Digital Health Hub and the EPSRC.
The authors have no competing interests to declare that are relevant to the content of this article. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the EU or the ERC. Neither the EU nor the granting authority can be held responsible for them.

### v3

@@ 1 Introduction
Current approaches for cerebrovascular analysis can be classified into three categories. Segmentation-based methods, enabled by recent advances in deep learning for accurate automated vessel segmentation, typically from time-of-flight magnetic resonance angiography (TOF-MRA), estimate arterial volume from the segmented masks at the whole-brain and regional levels [2]. Atlas-based approaches provide population references through statistical arterial atlases [3], offering normative cerebrovascular anatomy baselines. However, both approaches focus on simple volumetric and diameter measurements that may fail to completely characterize the vascular tree. Finally, graph-based computational tools use a segmentation mask as input to generate a graph representation of the vascular tree, from which intensity and morphometric features are extracted [4, 5, 6, 7, 8]. While commonly used, these tools are either semi-automatic [4, 7] or have been specifically designed for a particular type of image [5, 6, 8], limiting their usage at scale for any neurovascular imaging modality.
@@
Table 1: List of vascular features extracted within the vessel analysis framework.
@@
Abnormal-degree nodes $|\{v\in V:\deg(v)>3\}|$ Anomalous branching
@@
Fractal Fractal dimension $\frac{\Delta\log(N(\varepsilon))}{\Delta\log(\varepsilon)};$$N(\varepsilon)$: count of boxes at scale $\varepsilon$ Multi-scale organization
@@
Mean squared curvature $\frac{1}{L_{geo}}\int_{0}^{1}\frac{\kappa^{2}(t)}{[n(t)]^{2}}\|\mathbf{x}^{\prime}(t)\|dt$ Curvature variability
@@
In this work, we present CaravelMetrics, an automated framework for analyzing the cerebrovascular tree. Our framework extracts fifteen morphometric, topological, fractal, and geometric vascular features from a graph generated from vascular binary masks, thus designed to be agnostic to the underlying image modality. Similar to [2], we integrate an atlas of arterial regions, allowing us to analyze the cerebrovascular tree at both the whole-brain and regional levels. We showcase the potential of our open-source tool by analyzing the IXI Dataset [^1] https://brain-development.org/ixi-dataset/, a cohort of 570 subjects aged 20 to 86 years whose cerebrovascular trees were imaged using TOF-MRA. Our analysis reveals that our features effectively capture variations in the vessels across a population.
@@ 2 Methods
CaravelMetrics relies on representing the cerebrovascular network as a mathematical graph structure that is extracted from a vascular binary mask. The graph and the vascular binary mask are then used to extract morphometric, topological, fractal, and geometric features that can be aggregated at the whole-brain or regional level using an arterial atlas (Figure 1).
Graph Representation of Cerebrovascular Networks. Let us denote $M:\Omega\subset\mathbb{Z}^{3}\to\{0,1\}$, a binary vascular mask obtained from any 3D imaging modality. Given $M$, we define its baseline skeleton $S=\operatorname{Skel}(M)\subseteq\Omega$ as a one-voxel-thick subset that preserves the topological properties of $M$, such that $S\subseteq M$, $\operatorname{conn}(S)=\operatorname{conn}(M)$, and $\operatorname{hom}(S)=\operatorname{hom}(M)$, where $\operatorname{conn}(\cdot)$ denotes the number of connected components and $\operatorname{hom}(\cdot)$ describes the topology (loops and holes). From the skeleton points $\mathbf{x}=(x,y,z)\in S\subset\mathbb{Z}^{3}$, we construct an undirected weighted graph $G=(V,E)$, where each skeleton point corresponds to a node $v\in V$ with spatial coordinates $\mathbf{x}_{v}\in\mathbb{Z}^{3}$. Edges $E=\{(u,v)\}$ connect pairs of nodes within a $3\times 3\times 3$ neighborhood (26-connectivity).
After graph construction, two post-processing steps are applied. First, spurious edges are removed by sampling intensity along each edge and retaining low-intensity edges if $d_{geo}<2d_{euc}$, with $d_{geo}$ and $d_{euc}$ the geodesic and Euclidean distances between two edge endpoints. Second, to maximize coverage, orphan points, $\mathbf{x}\in M\setminus S$ with $\min_{s\in S}d_{uv}<D$, are identified and integrated via centroid connections under geodesic and branching-angle constraints.
@@
Feature Extraction. The framework extracts 15 vascular features, grouped into 4 categories (Table 1), each capturing a distinct property of cerebrovascular organization. Morphometric features quantify network extent and capacity through total vessel length and volume. Topological features describe branching architecture and connectivity using bifurcation count, loop number, and component analysis. Fractal features characterize hierarchical self-similarity via box-counting and lacunarity [9]. Finally, geometric features quantify vessel tortuosity, reflecting arterial stiffening and age-related alterations, by capturing local vessel morphology and shape irregularity through cubic-spline parameterization of discrete centerline segments (Eq. 1) as continuous curves $\mathbf{x}(t):[0,1]\!\to\!\mathbb{R}^{3}$, with curvature values weighted by path multiplicity $n(t)$ and averaged across root nodes $s\in\{R_{1},R_{2},R_{3}\}$ to obtain descriptors for each $G_{k}$.
@@
The cerebrovascular tree exhibits heterogeneous patterns across brain regions, motivating a territory-based analysis. To enable this, an arterial atlas $\mathcal{A}=\{T_{i}\}_{i=1}^{T}$, which partitions the brain into $T$ arterial territories, is aligned to the neurovascular image, yielding territory-specific masks $M_{i}=M\cap T_{i}$. Our framework supports two complementary analysis modes: whole-brain analysis, on the full graph $G=(V,E)$ derived from $M$; and region-specific analysis, on regional graphs $G_{i}=(V_{i},E_{i})$ derived from $M_{i}$. This formulation enables global or arterial-territory-specific feature extraction.
Implementation Details. Skeletonization, graph generation, and feature extraction are implemented in Python 3.10 with NumPy, SciPy, scikit-image, vedo, and NetworkX. Orphan connections are validated by a $30^{\circ}$ branching-angle constraint and a distance constraint $D>2.5$ mm. Graph geometry is refined via iterative Laplacian smoothing ($N_{iter}=2,\alpha=0.8$) and components with << 7 nodes are removed. We employ the open-source 3D brain MRI arterial territories atlas proposed in [10] ($T=30$). Atlas-to-image alignment is done through multi-step registration: brain extraction (FSL BET [11]), reorientation to MNI space, atlas-to-T1w registration using FLIRT with mutual information, and T1w-to-TOF-MRA co-registration. Code and documentation are available in the project’s repository [^2] https://github.com/i-vesseg/CaravelMetrics.
@@
3 Experiments and Results
We evaluate the proposed framework by analyzing 570 healthy subjects (20–86 years) from the IXI dataset, with multiple MR sequences acquired across three London hospitals (Guys, HH, and IOP) with different scanners. Each subject has associated demographic information, including age, sex, height, weight, education, and occupation.Body mass index (BMI) is computed from height and weight, and subjects are stratified into 4 standard BMI categories.
Using IXI’s TOF-MRA deep learning-generated binary vessel masks from the VesselVerse dataset [12], we apply CaravelMetrics to quantify age-related cerebrovascular changes and demographic associations with vessel morphology, tortuosity, and network topology, consistent with the literature (e.g., [13, 14]). Atlas-based regional analysis uses T1w IXI images for multi-step registration. Age effects are assessed using Spearman correlations ($r$) and one-way ANOVA across quartile-based groups with $\eta^{2}$ effect sizes. Demographic comparisons utilize t-tests for binary factors (sex) and one-way ANOVA for categorical variables (BMI, education), with $\omega^{2}$ effect sizes for unbalanced groups. Multi-center effects undergo site-stratified analysis to separate biological variation from acquisition artifacts. All analyses apply Benjamini-Hochberg False Discovery Rate (FDR) correction ($q<0.05$) to control for multiple comparisons.
@@ 3 Experiments and Results
Results. Initial analysis revealed substantial site effects, particularly in curvature, total length, and volume ($\omega^{2}=0.723$, $0.406$, $0.289$), with IOP showing markedly different vessel distributions than Guys and HH (Figure 2), likely reflecting scanner/acquisition variations [3], though age imbalances may have contributed. Excluding IOP and focusing on the two homogeneous sites ($n=488$) strengthened age effects: total vessel length effect size increased from $\eta^{2}=0.095$ to $0.274$ (17.9 percentage points), suggesting site heterogeneity attenuates biological associations. Analyzing age-related changes, negative correlations emerged between age and structural metrics (Fig. 3). Total length, bifurcation count, and fractal dimension showed the sharpest declines ($r\approx-0.55$, $p<0.001$), with total length exhibiting a  20% reduction across the lifespan (Fig. 4). Decreases in volume ($r\approx-0.38$), loop number ($r\approx-0.50$), and bifurcation density ($r\approx-0.35$) reflect network simplification and a loss of branching complexity exceeding vessel shortening. Conversely, curvature metrics ($r\approx+0.20$) and lacunarity ($r\approx+0.50$) increased with age, suggesting the remaining structure becomes more sparse and tortuous. Quartile-based age analysis revealed large effect sizes for total length ($\eta^{2}=0.274$), bifurcations ($\eta^{2}=0.276$), volume ($\eta^{2}=0.133$), lacunarity ($\eta^{2}=0.235$), fractal dimension ($\eta^{2}=0.220$), and bifurcation density ($\eta^{2}=0.132$) (all $p<0.001$), with a more modest association for mean curvature ($\eta^{2}=0.036$), confirming these features effectively capture structural vascular aging. Regional analysis (Fig. 5) revealed differences in correlation with age ($|r|\leq 0.30$) across cortical territories, with declines in length and fractal dimension with increased lacunarity, and deep nuclear regions exhibiting the opposite behavior.
@@
Regarding anthropometry, normal-weight individuals ($n=252$) had longer vessels than overweight ($n=146$) or obese ($n=55$) groups ($\eta^{2}=0.060$, $p<0.001$; Fig. 6, left). Higher BMI was associated with structural simplification, shown by reduced bifurcation and node counts ($\eta^{2}\approx 0.060$), increased curvature ($\eta^{2}=0.026$), and lacunarity ($\eta^{2}=0.059$). The underweight group ($n=3$) showed a similar pattern, though small sample size and age confounding limit the generalizability of the observations. Height showed a significant positive association with vascular volume ($\eta^{2}=0.037,p=0.007$). This is consistent with allometric scaling, where larger body size requires increased cerebral blood volume to perfuse a proportionally larger brain mass, without necessarily altering network complexity or density. Post-IOP removal, females ($n=265$) exhibited higher fractal dimension ($d=0.43$) and bifurcation density ($d=0.31$) than males ($n=223$), suggesting more compact, complex networks that mirror height-related allometric scaling ($d=0.39$ lacunarity in males). Education analysis was prioritized over occupation to minimize age confounding. Results revealed a clear stepwise gradient, with higher qualification levels associated with increased vessel length ($\omega^{2}=0.077$, $p<0.001$, $r=0.247$; Fig. 6, right), increased fractal dimension ($\omega^{2}=0.045$) and reduced lacunarity ($\omega^{2}=0.043$, suggesting a more complex, spatially homogeneous organization consistent with cognitive reserve extending to cerebrovascular architecture [15].
@@ 4 Discussion and Conclusion
We introduced CaravelMetrics, an automated framework for cerebrovascular characterization that extracts 15 morphometric, topological, fractal, and geometric features directly from binary vessel masks, making it compatible with diverse neurovascular imaging modalities. Atlas integration enables both whole-brain and territory-specific analyses, enabling multiscale assessment of cerebrovascular organization. Applied to 570 IXI subjects, CaravelMetrics revealed age-related reductions in vessel length, volume, bifurcation count, and fractal dimension, alongside increased tortuosity and lacunarity, as well as education- and sex-associated network differences, consistent with literature regarding length reductions and increased tortuosity in healthy aging [13, 14]. Limitations include sensitivity to mask quality and skeletonization. Future work should evaluate robustness across diverse protocols and pathological cohorts to validate these findings beyond a single healthy population.
@@ 6 Acknowledgments
This work is co-funded by the European Union (ERC CARAVEL 101171357) and France Life Imaging (ANR-11-INBS-0006). LSC is supported by TwinsUK Imaging: A Resource for Ageing Research (Chronic Disease Research Foundation). MM and SO are supported by King’s Health Partners Digital Health Hub and the EPSRC. The authors have no competing interests to declare. Views and opinions expressed are those of the authors and do not necessarily reflect those of the EU or the ERC. Neither the EU nor the granting authority can be held responsible for them. The authors thank Bogdan Ion and Duccio Lalli for their feedback on the graph representation.